# Decoding Time Series with LLMs: A Multi-Agent Framework for Cross-Domain Annotation

### v1

@@ 4.3 Evaluating General Annotations in Synthetic Datasets
We construct a synthetic dataset with time series data and ground-truth annotations to validate TESSA’s performance. Implementation details are provided in Appendix G.1.
Evaluation Metrics. We apply the LLM-as-a-judge approach Bubeck et al. (2023); Dubois et al. (2024), evaluating two metrics: Clarity and Comprehensiveness. Two distinct LLMs score the generated annotations on a scale of $1$ to $5$ for each metric, with an overall score calculated as the mean of the two metrics.
Experimental Results. We compare TESSA and DirectLLM in Table 2. The “Mean” represents the average score of generated annotations for each method, and P(T>D) is the percentage of TESSA’s annotations that receive higher scores than DirectLLM’s. The results show that TESSA outperforms DirectLLM across both metrics, with average scores of $3.90$ in Clarity and $4.44$ in Comprehensiveness, compared to DirectLLM’s $3.79$ and $1.55$. Additionally, $82.71\%$ of TESSA’s annotations receive higher scores, indicating that TESSA produces more essential and easily understandable features, further demonstrating its effectiveness.
@@ 4.4 Domain-specific Annotation Evaluation
In this subsection, we evaluate the quality of domain specific annotations. Similar to Section 4.3, we adopt a LLM-as-a-Judger strategy to evaluate the performance of domain-specific annotation agent from three perspectives: Clarity, Comprehensiveness, and Domain-relevance. The overall score is the average of these three metrics. Further details on these metrics are provided in Appendix H.1.
Experimental Results. We present the comparison results of TESSA and DirectLLM on the Environment dataset in Table 3, with GPT-4o as the LLM backbone. The key observations are: (1) TESSA significantly outperforms DirectLLM across all metrics, achieving an overall score of $4.64$ compared to DirectLLM’s $3.41$. Notably, $98.51\%$ of TESSA’s annotations receive higher scores, demonstrating its effectiveness in generating high-quality domain-specific annotations. (2) TESSA scores $4.74$ in Clarity and $4.38$ in Comprehensiveness, while DirectLLM scores $3.32$ and $3.01$, respectively. This shows that TESSA ’s annotations are clearer, more concise, and cover more important features. (3) TESSA also excels in domain relevance, with $94.72\%$ of its annotations scoring higher, achieving an average of $4.30$, significantly outperforming DirectLLM’s $3.41$. This indicates that TESSA produces highly accurate annotations that effectively use domain-specific terminology and maintain strong contextual relevance. Further results on other datasets are available in Appendix H.2.
Table 3: Domain-specific annotation results on the Environment dataset with GPT-4o as the LLM backbone.
@@ 4.5 In-depth Dissection of TESSA
Adaptive Feature Selections. We compare our two feature selection methods: offline LLM-based selection and incremental RL-based selection. To assess their effectiveness in selecting the top-$k$ most important features, we evaluate the quality of the generated general and domain-specific annotations, following the procedures in Sections 4.2 and 4.4, respectively. Environment is set as the target domain, with results shown in Fig. 2. The results indicate that TESSA performs comparably in both general and domain-specific annotation generation using either selection method. Specifically, as shown in Fig. 2(a), both approaches achieve MSE and MAE around $0.46$ and $0.44$ for general annotations. Similarly, in Fig. 2(b), both methods score consistently high across all domain-specific metrics, demonstrating their effectiveness in selecting important features. However, incremental RL-based selection proves more cost-effective by reducing redundant re-querying of previously used data.
@@
To prove the importance of adaptive feature selection in TESSA, we remove the adaptive feature selection module to create a variant, TESSA/F. We apply an LLM-as-a-judger to compare the quality of the generated annotations between TESSA and its variants. The evaluation metrics are introduced in Appendix I.1. We choose Social Good as the target dataset. The comparison results are presented in Table 4, with qualitative examples provided in Appendix I.2. We observe that: TESSA consistently outperforms TESSA/F. Specifically, TESSA achieves a clarity score of $4.41$, compared to $3.66$ for TESSA/F. This demonstrates the necessity of adaptive feature selection. Furthermore, according to Table 13 in Appendix I.2, the annotations generated by TESSA/F tend to include many features without proper analysis. This shows that involving too many features can hinder the clarity of the annotations, further emphasizing the importance of adaptive feature selection in improving annotation quality.
@@ 4.6 Case Study of TESSA
We conduct a case study to further validate the effectiveness of TESSA. A representative time series from the Social Good domain (Fig. 3(b)(b)) is selected, and both TESSA and DirectLLM are applied to generate general and domain-specific annotations, summarized in Table 14. To assess the quality of the annotations, we use an LLM-as-a-judger to evaluate the domain-specific annotations from both methods, with results shown in Table 12. Our findings indicate that: (1) TESSA’s general annotations capture more meaningful patterns, aiding user understanding and downstream tasks, whereas DirectLLM only highlights basic trends; and (2) TESSA’s domain-specific annotations consistently outperform DirectLLM across all metrics, offering clearer, more comprehensive, and contextually relevant insights. More case studies of multivariate time series data are provided in Appendix J.
@@ 5 Conclusion
Appendix A More Related Works
@@ D.1 Offline LLM-based Feature Selection
Appendix E Experimental Settings
E.1 Dataset Statistics
Datasets. To evaluate the effectiveness of TESSA, five real-world datasets from distinct domains are considered: Stock, Health, Energy, Environment, and Social Good. Specifically, the stock dataset includes 1,935 US stocks with the recent 6-year data. The other four datasets come from the public benchmark Time-MMD Liu et al. (2024a). The dataset statistics are summarized in Table 6.
@@ E.1 Dataset Statistics
E.2 Baseline Methods
@@
Appendix F Additional Results for General Annotation Evaluation in Downstream Tasks
F.1 Implementation Details
@@
F.2 Evaluation in Time Series Forecasting Tasks
@@
F.3 Evaluation in Time Series Imputation Tasks
@@
Implementation Details. We implement the multi-modal time series imputation based on TSLib Wu et al. (2023). We use Informer Zhou et al. (2021) as the forecasting model for the time series forecasting task. The model is configured with a dropout rate of $0.1$ and a learning rate of $0.0001$. GPT-4o is set as the LLM backbone. Other settings follow these in Section F.1.
@@
Appendix G Additional Details of General Annotation Evaluation in Synthetic Datasets
G.1 Implementation Details
@@
Appendix H Additional Results of Domain-specific Annotation Evaluation
H.1 Evaluation Metrics
@@
H.2 Additional Results on Other LLM Backbones
@@
Appendix I Additional Details of Ablation Studies
I.1 Evaluation Metric
To evaluate the effectiveness of adaptive feature selection, we use an LLM-as-a-judger to assess the general annotations generated by TESSA and its variant, TESSA/F.
I.2 Qualitative Examples
@@
Table 12: Case study: Evaluation results of domain-specific annotation of time series data in Figure 3(b) from the Environment dataset. GPT-4o is the LLM backbone.
@@
Table 13: Ablation studies of the impact of adaptive feature selection. Two compared methods are used for generating general annotation ploted in Figure 3(b)(a). Purple highlights basic patterns, green highlights general patterns.
@@
Appendix J Additional Details of Case Studies
@@
Case Study for Multivariate Time Series We then conduct a case study to demonstrate the effectiveness of TESSA in generating high-quality annotations for multivariate time series data. Specifically, we set the Finance dataset as the target domain. Health and Environment datasets are then applied in the source domains. The example multivariate time series data is shown in Fig. 5(d), where the multivarate time series data has four variables, i.e., price, volumn, relative strength index (RSI) and simple moving average (SMA). The generated annotations are shown in Table 15. From the table, we observe that (i) TESSA’s generated annotations are more naturally than DirectLLM; (ii) DirectLLM interprets each variable independently by only focusing their trends. However, TESSA can capture the correlation between variables. This shows TESSA is able to analyze inter-variable patterns. These further imply the effectiveness of TESSA in generating high-quality domain-specific annotations for multivariate time series data.
@@
Additional Examples on Various Domains. Additional examples are presented for the synthetic, environment, energy and social good datasets, respectively. Specifically, the general annotations of selected time series on the synthetic dataset that in Fig. 6 are shown in Tables 17 and 18. The domain-specific annotations of selected time series on the environment dataset (Fig. 7) are shown in Tables 19 and 20. And the domain-specific annotations of time series on the energy dataset (Fig. 8) are shown in Tables 21 and 22. Similarly, the domain-specific annotations of time series on the social good dataset (Fig. 9) are shown in Tables 23 and 24. Similar observations to those in Table 12 and Table 16 are found.
@@
Table 14: A case study in Social Good dataset. The target time series data is shown in Figure 3(b) (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights domain-specific patterns.

### v2

@@ 4.3 Evaluating General Annotations in Synthetic Datasets
We construct a synthetic dataset with time series data and ground-truth annotations to validate TESSA’s performance. Implementation details are provided in Appendix H.1.
Evaluation Metrics. We apply the LLM-as-a-judge approach Bubeck et al. (2023); Dubois et al. (2024), evaluating two metrics: Clarity and Comprehensiveness. Two distinct LLMs score the generated annotations on a scale of $1$ to $5$ for each metric, with an overall score calculated as the mean of the two metrics. Further details on the metrics and the LLM-judge prompts can be found in Appendix H.1.
Experimental Results. We compare TESSA with DirectLLM in Table 2. The “Mean” denotes the average score of generated annotations for each method, and P(T>D) is the percentage of TESSA’s annotations that receive higher scores than DirectLLM’s. The results show that TESSA outperforms DirectLLM on both metrics, with average scores of $3.90$ in Clarity and $4.44$ in Comprehensiveness, compared to DirectLLM’s $3.79$ and $1.55$. Additionally, $82.71\%$ of TESSA’s annotations receive higher scores, indicating that TESSA produces more essential and easily understandable features, further demonstrating its effectiveness.
@@ 4.4 Domain-specific Annotation Evaluation
In this subsection, we evaluate the quality of domain specific annotations. Similar to Section 4.3, we adopt a LLM-as-a-Judger strategy to evaluate the performance of domain-specific annotation agent from three perspectives: Clarity, Comprehensiveness, and Domain-relevance. The overall score is the average of these three metrics. Further details on these metrics are provided in Appendix I.1.
Experimental Results. We present the comparison results of TESSA and DirectLLM on the Environment dataset in Table 3, with GPT-4o as the LLM backbone. The key observations are: (1) TESSA significantly outperforms DirectLLM across all metrics, achieving an overall score of $4.64$ compared to DirectLLM’s $3.41$. Notably, $98.51\%$ of TESSA’s annotations receive higher scores, demonstrating its effectiveness in generating high-quality domain-specific annotations. (2) TESSA scores $4.74$ in Clarity and $4.38$ in Comprehensiveness, while DirectLLM scores $3.32$ and $3.01$, respectively. This shows that TESSA ’s annotations are clearer, more concise, and cover more important features. (3) TESSA also excels in domain relevance, with $94.72\%$ of its annotations scoring higher, achieving an average of $4.30$, significantly outperforming DirectLLM’s $3.41$. This indicates that TESSA produces highly accurate annotations that effectively use domain-specific terminology and maintain strong contextual relevance. More results on other datasets are in Appendix I.2.
Table 3: Domain-specific annotation results on the Environment dataset using GPT-4o as the LLM. Dom. Rel. is the domain-relevance metric used in Section 4.4.
@@ 4.5 In-depth Dissection of TESSA
Adaptive Feature Selections. We compare our two feature selection methods: offline LLM-based selection and incremental RL-based selection. To assess their effectiveness in selecting the top-$k$ most important features, we evaluate the quality of the generated general and domain-specific annotations, following the procedures in Sections 4.2 and 4.4. Environment is set as the target domain, with results shown in Fig. 3. The results indicate that TESSA performs comparably in both general and domain-specific annotation generation using either selection method. Specifically, as shown in Fig. 3(a), both approaches achieve MSE and MAE around $0.46$ and $0.44$ for general annotations. Similarly, in Fig. 3(b), both methods score consistently high across all domain-specific metrics, demonstrating their effectiveness in selecting important features. However, incremental RL-based selection proves more cost-effective by reducing redundant re-querying of previously used data.
@@
To prove the importance of adaptive feature selection in TESSA, we remove the adaptive feature selection module to create a variant, TESSA/F. We apply an LLM-as-a-judger to compare the quality of the generated annotations between TESSA and its variants. The evaluation metrics are introduced in Appendix L.1. We choose Social Good as the target dataset. The comparison results are presented in Table 4, with qualitative examples provided in Appendix L.2. We observe that: TESSA consistently outperforms TESSA/F. Specifically, TESSA achieves a clarity score of $4.41$, compared to $3.66$ for TESSA/F. This demonstrates the necessity of adaptive feature selection. Furthermore, according to Table 15 in Appendix L.2, the annotations generated by TESSA/F tend to include many features without proper analysis. This shows that involving too many features can hinder the clarity of the annotations, further emphasizing the importance of adaptive feature selection in improving annotation quality. Additional ablation studies examining the contributions of other components of TESSA are in Appendix L.3. And discussions on data contamination are provided in Appendix L.4.
@@ 4.6 Case Study of TESSA
We conduct a case study to further validate the effectiveness of TESSA. A representative time series from the Social Good domain (Fig. 6(b)(b)) is selected, and both TESSA and DirectLLM are applied to generate general and domain-specific annotations, summarized in Table 22. To assess the quality of the annotations, we use an LLM-as-a-judger to evaluate the domain-specific annotations from both methods, with results shown in Table 14. Our findings indicate that: (1) TESSA’s general annotations capture more meaningful patterns, aiding user understanding and downstream tasks, whereas DirectLLM only highlights basic trends; and (2) TESSA’s domain-specific annotations consistently outperform DirectLLM across all metrics, offering clearer, more comprehensive, and contextually relevant insights. More case studies of multivariate time series data are provided in Appendix M.
@@ 5 Conclusion
6 Limitations
Potential limitations of this work include the need for limited target-domain annotations to learn domain-specific jargon for generating domain-specific annotations. Additionally, our approach relies on annotations from source domains to transfer knowledge to target domains. If the chosen source domain annotations are all of low quality or lack sufficient common knowledge, it may affect the overall performance of TESSA.
7 Ethics Statement
We adhere to the ACM Code of Ethics in our research. All datasets and models used in this study are either publicly accessible or synthetically generated. Specifically, we created a synthetic dataset comprising time series data with generated general annotations to facilitate our experiments while avoiding the use of any personal or sensitive real-world data. We acknowledge the potential risks and harms associated with LLMs, such as generating harmful, offensive, or biased content. Moreover, LLMs are often prone to generating incorrect information, sometimes referred to as hallucinations. We recognize that the models studied in this paper are not exceptions to these limitations. Previous research has shown that the LLMs used in this study suffer from bias, hallucinations, and other issues. We emphasize the importance of responsible and ethical use of LLMs and the need for further research to mitigate these challenges before deploying them in real-world applications. The models used in this work are licensed under the terms of OpenAI, LLaMA, and Qwen.
Appendix A More Related Work
@@ Appendix B Notations
Additionally, we also provide some specific examples of domains, annotations, and features to improve the clarity of the problem settings of cross-domain time series annotation defined in Section 3. Specifically, our paper consider six distinct domains, i.e., Stock, Health, Environment, Social Good, Climate and Economy. For instance, the stock dataset includes multivariate time series data (e.g., stock price, volume, RSI, moving average) with corresponding annotations capturing features like support levels and resilience. More examples of text-wise features can be found in Table 13. More examples of domain-specific annotations are provided in Tables 2728293031 and 32 of Appendix M.‘
@@ D.1 Offline LLM-based Feature Selection
D.2 Incremental Reinforcement Learning-based Feature Selection
The necessity of this component. In the proposed LLM-based feature selection from Section 3, when new annotations exhibit different distributions or feature characteristics compared to the old data, it becomes necessary to re-query both old and new data to select the top-$k$ most important features. This process is computationally intensive and resource-inefficient, especially as the volume of data grows. To address this issue, we propose the incremental reinforcement learning (RL)-based feature selection method. This approach provides the following benefits:
  • Cost-Efficiency: Instead of re-querying LLMs with all the data, the RL-based method incrementally updates the knowledge stored in policy networks, requiring only the new data during updates.
  • Scalability: By reducing redundant computations, the incremental RL-based method ensures scalability in dynamic environments with evolving data.
Appendix E Time Complexity Analysis
To analyze the time complexity of TESSA, we consider each component of TESSA separately, focusing on the computational cost associated with feature extraction, feature selection, and annotation generation and review.
Feature Extraction. Extracting intra-variable and inter-variable features has a complexity of $\mathcal{O}(C^{2}\cdot n)$ , where $C$ is the number of channels in the time series and $n$ is the number of time points.
Feature Selection. For offline LLM-based feature selection, the complexity is $\mathcal{O}(k\cdot M\cdot L^{2})$, where $k$ is the number of features, $M$ is the model size (number of parameters) of the LLM, and $L$ is the input sequence length. Incremental RL-based selection reduces this overhead by incrementally updating policy networks without re-querying old data.
Annotation Generation and Review. Each LLM inference for annotation generation or review has a complexity of $\mathcal{O}(M\cdot L^{2})$. Given $T$ samples, the overall complexity becomes $\mathcal{O}(T\cdot M\cdot L^{2})$.
Overall Complexity. The combined complexity can be expressed as:
$\mathcal{O}(T\cdot[C^{2}\cdot n+k\cdot M\cdot L^{2}]),$
where $T$ is the number of time series samples.
Appendix F Experimental Settings
F.1 Dataset Statistics
Datasets. To evaluate the effectiveness of TESSA, five real-world datasets from distinct domains are considered: Stock, Health, Energy, Environment, and Social Good. Specifically, the stock dataset includes 1,935 US stocks with the recent 6-year data, collected from Investtech [^2] https://www.investtech.com/. The other four datasets come from the public benchmark Time-MMD Liu et al. (2024a). The dataset statistics are summarized in Table 6.
@@ F.1 Dataset Statistics
Climate Monthly 5 496 177
Economy Monthly 3 423 410
F.2 Baseline Methods
@@
F.3 Framework for Multi-modal Downstream tasks
To evaluate the quality of general annotations, we leverage the multi-modal time series analysis framework proposed in Time-MMD Liu et al. (2024a), illustrated in Fig. 4. Using time series forecasting as a representative task, this framework employs an end-to-end pipeline that combines open-source language models with diverse time-series forecasting (TSF) models. Time-series data and textual annotations are modeled independently through dedicated unimodal TSF architectures and language models (LLMs) equipped with projection layers. The outputs of these modalities are fused via a dynamic linear weighting mechanism to generate final predictions. To optimize computational efficiency, we keep the LLM parameters frozen during training and update only the projection layers. Additionally, pooling layers are introduced to resolve dimension mismatches between textual variables and time-series features. The framework supports end-to-end training with minimal parameter overhead, ensuring both scalability and practicality.
Figure 4: Overall framework of MM-TSFlib from Time-MMD Liu et al. (2024a) used in our multi-modal downstream tasks. MMTSFlib uses a model-agnostic multimodal integration framework that independently models time-series and textual annotations within an end-to-end training manner. MM-TSFlib slightly increases the number of trainable parameters, balancing effectiveness and efficiency.
Appendix G Additional Results for General Annotation Evaluation in Downstream Tasks
G.1 Implementation Details
@@
G.2 Evaluation in Time Series Forecasting Tasks
@@
G.3 Evaluation in Time Series Imputation Tasks
@@
Implementation Details. We implement the multi-modal time series imputation based on TSLib Wu et al. (2023). We use Informer Zhou et al. (2021) as the forecasting model for the time series forecasting task. The model is configured with a dropout rate of $0.1$ and a learning rate of $0.0001$. GPT-4o is set as the LLM backbone. Other settings follow these in Section G.1.
@@
Appendix H Additional Details of General Annotation Evaluation in Synthetic Datasets
H.1 Implementation Details
@@
Appendix I Additional Results of Domain-specific Annotation Evaluation
I.1 Evaluation Metrics
@@
I.2 Additional Results on Other LLM Backbones
@@
Appendix J Experimental results of Human in Loop
In this section, we demonstrate that the annotations generated by TESSA can assist humans in analyzing time-series data. To evaluate this, we selected $60$ time-series samples from three domains—Environment, Energy, and Social Good—with $20$ datasets per domain. We compared annotations generated by TESSA and DirectLLM, asking $20$ PhD stduents, researchers, and professors as the participants to assess which annotations were more informative and useful. The results of this human-in-the-loop evaluation, summarized in Table 11, reveal that TESSA consistently outperforms DirectLLM across all three domains. Participant assessments indicate that $88.3\%$ of TESSA’s general annotations and $93.3\%$ of its domain-specific annotations are more informative compared to DirectLLM’s outputs. This substantiates TESSA ’s capacity to produce semantically meaningful annotations that enhance human interpretability during time-series analysis workflows.
Table 11: Comparison results of general and domain-specific annotations. GPT-4o is the LLM backbone.
Domain Method P(T>D) (%)
Environment General 83.3
Specific 91.7
Energy General 88.3
Specific 93.3
Social Good General 85.8
Specific 92.5
Appendix K Additional Results on Multivariate Time Series
In this section, we aim to demonstrate the effectiveness of TESSA in generalizing to multivariate time series. Specifically, We use Stock and Health datasets as the source domains, Climate and Economy datasets as the target domains.
Domain-specific Annotations Evaluation. Similar to Section 4.4, we adopt a LLM-as-a-Judger strategy to evaluate the performance of TESSA and DirectLLM in generating domain-specific annotations. Other settings follow these in Section 4.4. We present the comparison results on the two datasets in Table 12. From the table, we observe that similar to these of univariate time series in Section 4.4, TESSA significantly outperforms DirectLLM across all metrics in both two multivariate time series datasets, aciving an overall score of $4.51$ and $4.55$ compared to DirectLLM’s $3.38$ and $3.42$ in Climate and Economy dataset, respectively. This demonstrates the effectiveness of TESSA in generalizing to multivariate time series. Additional case studies of TESSA applied to multivariate time series are presented in Appendix M.
Table 12: Comparison results of domain-specific annotations for multivariate time series in the Climate and Economy datasets using GPT-4o as the LLM backbone.
Domains Metric Method Mean P(T>D) (%)
Climate Clarity TESSA 4.58 99.17
DirectLLM 3.35
Compre. TESSA 4.37 96.43
DirectLLM 3.38
Dom. Rel. TESSA 4.57 96.20
DirectLLM 3.40
Overall TESSA 4.51 97.26
DirectLLM 3.38
Economy Clarity TESSA 4.46 97.14
DirectLLM 3.48
Compre. TESSA 4.57 96.81
DirectLLM 3.41
Dom. Rel. TESSA 4.61 96.70
DirectLLM 3.37
Overall TESSA 4.55 96.88
DirectLLM 3.42
Appendix L Additional Details of Ablation Studies
L.1 Evaluation Metric
To evaluate the effectiveness of the adaptive feature selection, we use an LLM-as-a-judger to evaluate the general annotations generated by TESSA and its variant TESSA/F.
L.2 Qualitative Examples
@@
Table 14: Case study: Evaluation results of domain-specific annotation of time series data in Fig. 6(b) from the Environment dataset. GPT-4o is the LLM backbone.
@@
Table 15: Ablation studies of the impact of adaptive feature selection. Two compared methods are used for generating general annotation ploted in Fig. 6(b)(a). Purple highlights basic patterns, green highlights general patterns.
@@
L.3 Additional Ablation Studies on the Contributions of TESSA’s Components
In Section 4.5, we conduct ablation studies to assess the importance of domain decontextualization and adaptive feature selection in TESSA. To further understand the contributions of TESSA’s components, we conduct additional ablation studies on domain-specific term extractor and annotation reviewers, respectively.
Domain-specific Term Extractor. To understand the ability of LLMs to process domain-specific jargon to generate useful features, we implement a variant TESSA/S, which removes the domain-specific term extractor in TESSA and directly generate domains-specific annotations from general annotations. We provide the comparison results in Stock and SocialGood datasets in Tables 16 and 17, which demonstrate that TESSA is able to capture more jargon-rich in information, such as resistance levels in the Stock dataset and reactive bounds in the SocialGood dataset. In contrast, TESSA/S merely converts general features into basic domain-specific features. This highlights that LLMs, when used as domain-specific term extractors, can effectively generate valuable jargon that enhances the generation of domain-specific annotations.
Annotation Reviewer. To evaluate the impact of the annotation reviewer component, we introduce a variant TESSA/A that excludes this module. We assess the quality of domain-specific annotations generated by TESSA and TESSA/A using an LLM-as-a-judger framework, focusing on two criteria: (1) Clarity (readability and coherence of annotations) and (2) Domain-Relevance (alignment with domain-specific context). An overall score, calculated as the average of these metrics, provides a holistic performance measure. As shown in Table 18, TESSA achieves statistically significant improvements over TESSA/A across all metrics. This quantitative comparison demonstrates that the annotation reviewer critically enhances both the clarity and contextual relevance of generated annotations, ensuring they better capture domain-specific nuances.
Table 16: Ablation studies of the impact of the domain-specific term extractor in Stock dataset. Purple highlights basic patterns, green highlights general patterns, yellow highlights domain-specific patterns and blue highlights correlations between variables in multivariate time series data.
TESSA: Compass Digital Acquisition Unit’s stock price shows a notable pattern of rising and falling periodically, indicating seasonality with stable long-term trends interrupted by short-term fluctuations. There are key resistance levels around intervals 134,270, and 403, where prices peak before dipping. The stock volume demonstrates significant spikes at specific points, suggesting irregular activity, particularly around values 7,000 and 80,500, which may indicate volume bursts or unusual market events. The relative strength index (RSI) also reveals a recurring pattern, gradually trending upward before a sharp decline, reflecting a cycle of growth and subsequent drop. Overall, the mild positive correlation between stock price and RSI indicates that periodic changes in price are somewhat echoed in RSI patterns, potentially offering predictive insights for future stock movements.
TEESA/S: Compass Digital Acquisition Unit’s stock price exhibits a distinct cyclical pattern, suggesting seasonality, with occasional fluctuations. The relative strength index (RSI) shows a repeating trend, gradually rising before experiencing a sharp drop, indicating a pattern of growth followed by a decline. The mild positive relationship between stock price and RSI suggests that changes in price tend to be mirrored in the RSI, offering potential clues for predicting future price movements.
Table 17: Ablation studies of the impact of the domain-specific term extractor in SocialGood dataset. Purple highlights basic patterns, green highlights general patterns and yellow highlights domain-specific patterns.
TESSA: The unemployment rate displays a clear upward trend over time, starting from the mid-5s and progressing to the mid-7s toward the end of the series, signaling persistent growth in unemployment. This movement is influenced by volatility periods, suggesting that economic cycle phases are periodically impacting employment levels. The observed resistance thresholds around the mid-5s and mid-6s mark critical threshold levels, where unemployment temporarily stabilizes before continuing its upward momentum. Analyzing lag features, past values like 5.5 and 6.0 serve as baselines, helping to understand how unemployment has evolved over time. This steady upward movement, punctuated by cyclical variations, highlights the need for strategies that address both economic momentum indicators and short-term reactive bounds, ensuring a robust response to both long-term trends and periodic disruptions in the labor market.
TEESA/S: The unemployment rate data reveals a consistent upward trend over time, starting from 5.18 and rising to approximately 7.24. This suggests an increasing trend in unemployment, which might pose socio-economic challenges if it persists. The data also exhibits seasonal fluctuations, indicating periods of elevated unemployment that could align with specific economic or legislative cycles. These recurring ups and downs highlight how external factors might periodically impact the job market. Recognizing these cyclical variations can help policymakers craft effective strategies to mitigate the potential socio-economic risks of rising unemployment in the future.
Table 18: Ablation studies for the annotation reviewer in the Social Good dataset with GPT-4o as the LLM backbone.
Metric Method P(TESSA>TESSA/A) (%)
Clarity TESSA 95.4
TESSA/A
Dom. Rel. TESSA 87.6
TESSA/A
Overall TESSA 91.6
TESSA/A
Source Type
UTF-8
bytes
Docs
Tokens
Common Crawl web pages 9,812 3,734 2,180
GitHub code 1,043 210 342
Reddit social media 339 377 80
Semantic Scholar papers 268 38.8 57
Project Gutenberg books 20.4 0.056 5.2
Wikipedia encyclopedic 16.2 6.2 3.7
Total 11,519 4,367 2,668
Table 19: The categority of the training data used to pre-train OLMo-7B Groeneveld et al. (2024), which is from OLMo’s technical report Groeneveld et al. (2024).
L.4 Additional Ablation Studies of Data Contamination
To investigate whether TESSA retains its effectiveness when the LLMs used are not exposed to data describing the time series, we conduct an ablation study using fully open LLMs, specifically OLMo-7B Groeneveld et al. (2024). The categories of the training data for the LLM are listed in Table 19. In this study, we carefully select time series with textual annotations that neither appear in the training data of OLMo-7B nor describe the time series themselves. We use the Stock and Health datasets as source domains, and the Energy dataset as the target domain. We then apply TESSA and DirectLLM to generate both general and domain-specific annotations for the time series data in the Energy dataset.
The results are presented in Table 20. From the table, we observe that: (1) despite using a fully open LLM that has not been exposed to data describing the time series, TESSA with OLMo-7B as the LLM backbone is still able to understand and generate informative annotations with richer general and domain-specific patterns, using more natural language. In contrast, DirectLLM only offers a simplistic description of the basic trend of the time series data. This further underscores the effectiveness of TESSA even in a scenario of strict data non-contamination.
Additionally, we employ an LLM-as-a-Judger strategy to evaluate the domain-specific annotations generated by TESSA and DirectLLM, as shown in Table 20. The evaluation is conducted from three perspectives: Clarity, Comprehensiveness, and Domain-relevance. GPT-4o serves as the LLM judger, and the other settings follow those in Sec. 4.4. The comparison results are presented in Table 21. We observe that TESSA consistently achieves full scores in all three metrics, significantly outperforming DirectLLM. This further demonstrates TESSA ’s effectiveness in generating high-quality domain-specific annotations with stronger clarity, comprehensiveness, and domain-relevant contextual precision.
Table 20: Ablation studies on data contamination using OLMo-7B-Instruct. We select a time-series in the Energy dataset, which is shown in Fig. 5. Purple highlights basic patterns, green highlights general patterns, and yellow highlights domain-specific patterns.
TESSA’s general annotations The time series data shows a trend of increasing and decreasing values with some seasonality, which can be observed from the values of the intra-variable time series features such as trend, seasonality, and moving average. The values of the intra-variable moving average are not stable, indicating that the time series data is volatile and prone to fluctuations. The values of the intra-variable lag features are not significant, indicating that there is no strong correlation among the time series data points. DirectLLM’s general annotations This time series data represents a process with high variability in the values, with an increasing trend. The data is separated by ’|’ tokens, which could represent different stages or conditions of the process. The highest value of 1.108 is observed at the 11th instance, which might indicate a bottleneck or peak in the process. The series shows a decreasing trend from instances 9-12, which suggests that certain measures or interventions were put in place to address the bottleneck. Overall, the data suggests that the process is subject to fluctuations, but can be managed by identifying and addressing the root causes of the bottlenecks.
TESSA’s domain-specific annotations The natural gas price is exhibiting a volatile pattern with significant fluctuations, suggesting a lack of stability in the market. This volatility hints at underlying market dynamics that are affecting price movements, potentially tied to factors such as supply and demand shifts or external economic variables. A closer examination of the trend channels indicates intermittent rising and falling trends, pointing towards a potential cyclical behavior in natural gas prices. Currently, the market does not show strong correlation from lag features, suggesting that recent price changes are not strongly influenced by past values. Support and resistance levels could play a crucial role, as breaking through these levels may signal a significant trend shift in the gas market’s future pricing. DirectLLM’s domain-specific annotations Gasonline prices in the Energy domain are fluctuating, ranging from $1.04$ to $1.10$ per unit, with occasional spikes above $1.10$, and drops below $1.00$. These prices fluctuate consecutively, indicating a dynamic market for Gasonline.
Figure 5: Selected time series data from Energy dataset for ablation studies on data contimination. This time series data has 30 data points. The corresponding generated annotations of TESSA and DirectLLMs are provided in Table 20.
Table 21: Ablation studies of data contamination. We use an LLM-as-a-Judger to compare a domain-specific annotation in the Energy dataset generated by DirectLLM and TESSA. OLMo-7B is the LLM-backbone. GPT-4o is used as the LLM judger.
Metric Method Mean
Clarity TESSA 5
DirectLLM 3
Compre. TESSA 5
DirectLLM 2
Dom. Rel. TESSA 5
DirectLLM 3
Overall TESSA 5
DirectLLM 2.67
Appendix M Additional Details of Case Studies
@@
Case Study for Multivariate Time Series We then conduct a case study to demonstrate the effectiveness of TESSA in generating high-quality annotations for multivariate time series data. Specifically, we set the Stock dataset as the target domain. Health and Environment datasets are then applied in the source domains. The example multivariate time series data is shown in Fig. 8(d), where the multivariate time series data has four variables, i.e., price, volume, relative strength index (RSI) and simple moving average (SMA). The generated annotations are shown in Table 23. From the table, we observe that (1) TESSA’s generated annotations are more natural than DirectLLM; (2) DirectLLM interprets each variable independently by only focusing their trends. However, TESSA can capture the correlation between variables. This shows TESSA is able to analyze inter-variable patterns. These further imply the effectiveness of TESSA in generating high-quality domain-specific annotations for multivariate time series data.
@@
Additional Examples on Various Domains. Additional examples are presented for the synthetic, environment, energy and social good datasets, respectively. Specifically, the general annotations of selected time series on the synthetic dataset that in Fig. 9 are shown in Tables 25 and 26. The domain-specific annotations of selected time series on the environment dataset (Fig. 10) are shown in Tables 27 and 28. And the domain-specific annotations of time series on the energy dataset (Fig. 11) are shown in Tables 29 and 30. Similarly, the domain-specific annotations of time series on the social good dataset (Fig. 12) are shown in Tables 31 and 32. Similar observations to those in Table 14 and Table 24 are found.
@@
Table 22: A case study in Social Good dataset. The target time series data is shown in Fig. 6(b) (a). Purple highlights basic patterns, green highlights general patterns, and yellow highlights domain-specific patterns.